at which time, 1:35 or 1:38, do you think opposite rays form?
1
Expert's answer
2017-08-22T14:31:07-0400
For 1 hour "hour ray" goes: 360/12=30 deg
For 1 min "minute ray" goes: 360/60=6 deg
Let's take 1:38 or 98 min,
than:
"hour ray" goes: 98*30/60=49 deg
"minute ray" goes: 98*6=588 deg or 588-360=228 deg
49+180 (when they form a straight line) =229 deg (approximately 228 deg).
If we do the same calculations for the time 1:35 we will get
1:38 or 95 min:
"hour ray" goes: 95*30/60=47.5 deg
"minute ray" goes: 95*6=570 deg or 570-360=210 deg
47.5+180 (when they form a straight line) =227.5 deg (that do not match 210 deg).
That's why 1:38 is a correct answer.